A wedge shaped film has refractive index 1.34 and thickness of extreme sides 0 (zero) and t. If a light of wavelength 492 nm is incident normally on it and 20 fringes are obtained, determine t.
1
Expert's answer
2020-04-05T14:49:39-0400

As per the given data in the question,

Refractive index of the film=1.34

thickness of the extreme side 0 and t

wavelength of the light "(\lambda)=492nm"

Number of fringes=20

As per the rule of x-ray reflection on the single layer

"2t\\sqrt{n^2\u2212\\sin^2i}-\\dfrac{\\lambda}{2}\u200b=m\u03bb"

n=refractive index of the medium

Light is getting incident normally, so "i=0"

So,

"2tn-\dfrac{\lambda}{2}=m\lambda"


"\Rightarrow t=\dfrac{m\lambda+\dfrac{\lambda}{2}}{2n}"


"\Rightarrow t=\dfrac{20\times492\times10^{-9}+246\times 10^{-9}}{2\times 1.34}"


"\Rightarrow t=3763\times10^{-9}m"


"t=3.76\mu m"